OpenStudy (anonymous):

csc x= (4√3)/3 and cos x<0, evaluate the function below: sin x

OpenStudy (ankit042):

@lamp12345 as already mentioned csc x = 1/sinx also as cosx is less than zero this means it is in second quadrant

OpenStudy (anonymous):

well if thats the case then 1*csc x would be obvious but thats not answer.

OpenStudy (ankit042):

sinx = 1/csc x not 1*csc x

OpenStudy (anonymous):

still not getting it.

OpenStudy (ankit042):

sinx = \[\sqrt{3}/4\]

sam (.sam.):

Look at it like this, \[\csc x= \frac{4\sqrt{3}}{3}\] We know that, from trig identity\[\csc{x}=\frac{1}{\sin{x}}\] So \[\frac{1}{\sin(x)}=\frac{4\sqrt{3}}{3} \\ \\ \sin(x)=\frac{3}{4\sqrt{3}}\]